# Gravitational quasinormal modes of Dymnikova black holes

### v1

@@ I Introduction
One of the most powerful tools of black-hole physics comes from the study of quasinormal modes (QNMs) Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Bolokhov and Skvortsova (2025a), which govern the damped oscillations of spacetime following a perturbation. These modes encode key information about the underlying geometry and dominate the ringdown phase of gravitational-wave signals detected by LIGO, Virgo, and KAGRA Abbott et al. (2016a, 2017, 2020, b). The spectrum of gravitational perturbations provides a unique fingerprint of the compact object, enabling tests of general relativity in the strong-field regime and offering the possibility of constraining or ruling out alternative models of gravity and exotic compact objects. For this reason, accurate determination of gravitational QNMs has become a central task in both theoretical and observational astrophysics.
@@
Among the various proposals, the solution introduced by Dymnikova Dymnikova (1992) stands out as an elegant example. It describes a static, spherically symmetric, asymptotically flat geometry in which the central singularity is replaced by a de Sitter core, leading to finite curvature invariants everywhere. While originally formulated as a phenomenological construction, the same functional form has more recently reappeared in the framework of Asymptotic Safety, where it can be derived from renormalization-group improvements tied to curvature invariants Platania (2019). Thus, the Dymnikova spacetime provides not only one of the earliest regular black-hole models, but also a concrete realization of quantum-gravity inspired corrections. This dual interpretation makes it a natural and timely candidate for detailed studies of gravitational perturbations and their observational signatures.
@@ II The Dymnikova Black Hole Geometry
The quest for singularity-free black-hole solutions has long motivated the construction of effective geometries that remain regular throughout spacetime. A particularly influential example was proposed by Dymnikova Dymnikova (1992), who suggested replacing the central singularity of the Schwarzschild solution with a de Sitter core while preserving asymptotic flatness. This is achieved through a stress–energy distribution that formally violates some of the classical energy conditions but can be regarded as an effective description of quantum backreaction. The resulting geometry is a static, spherically symmetric, and asymptotically flat spacetime in which all curvature invariants remain finite.
@@
describing a regular de Sitter interior, while for large $r$ one recovers the Schwarzschild behavior $f(r)\approx 1-2M/r$. The spacetime thus interpolates smoothly between two familiar limits while remaining free of singularities.
Although originally introduced as a phenomenological model, essentially the same functional form was later derived within the framework of Asymptotic Safety. In this approach, the Schwarzschild lapse function
@@
The Dymnikova solution enjoys a dual origin: historically as a phenomenological model of singularity resolution and, more recently, as a concrete realization of renormalization-group improvements in Asymptotic Safety. This dual perspective underscores its relevance as a minimal yet robust model of a regular black hole. Furthermore, higher-dimensional generalizations of the Dymnikova geometry have been shown to arise naturally in theories with higher-curvature corrections Konoplya and Zhidenko (2024a), further highlighting its role as a testing ground for semiclassical and quantum-gravity inspired modifications of black-hole physics.
@@ III Gravitational perturbations
Studying gravitational perturbations in spacetimes that incorporate quantum corrections is more subtle than in the classical case. The challenge arises because such geometries, including the Dymnikova black hole, are not derived as exact solutions of Einstein’s equations with a specified matter source, but rather emerge from effective constructions motivated by quantum-gravity considerations. As a result, a straightforward derivation of perturbation equations is not possible. A practical resolution was suggested in the loop-inspired framework of Ashtekar, Olmedo, and Singh Ashtekar et al. (2018a, b), where the background geometry can be reinterpreted within Einstein’s theory if one introduces an effective anisotropic fluid stress tensor that captures the relevant quantum corrections. This reformulation allows the machinery of black-hole perturbation theory to be applied in a consistent manner.
Following this idea, Bouhmadi-López and collaborators Bouhmadi-López et al. (2020); Konoplya and Stashko (2025a) analyzed axial perturbations under the assumption that fluctuations aligned with the anisotropy do not contribute in the axial sector. This simplification is reminiscent of the treatment in Bronnikov et al. (2012), where the decoupling of scalar and axial gravitational modes was exploited in scalar–tensor models (see also Chen and Chen (2019)). Although such assumptions necessarily restrict the analysis, they are widely adopted in related contexts and are generally expected to provide reliable results for small deviations from Schwarzschild geometry Konoplya and Stashko (2025b, a); Dubinsky (2025a); Bolokhov and Skvortsova (2025c).
@@
Figures 1 and 2 show the effective potentials for axial modes in the Dymnikova background. When the quantum parameter $l_{\rm cr}$ is very small (e.g., $l_{\rm cr}=0.01M$), the potential nearly coincides with the Schwarzschild one except in the immediate vicinity of the horizon. The deformation is highly localized, and the potential rapidly merges with the Schwarzschild profile at moderate radii. At the same time, the location of the event horizon shifts inward as $l_{\rm cr}$ grows, leading to a decrease in the horizon radius compared to the Schwarzschild value. These features indicate that the quantum-inspired modifications are concentrated in the near-horizon region, leaving the asymptotic structure essentially unaffected.
@@ IV Quasinormal modes
The WKB–Padé method is most reliable for moderate to large multipoles and low overtones (heuristically $n\ll\ell$ for single-barrier potentials). Tables I–IV summarize the results obtained with 6th and higher-order WKB methods combined with Padé resummation, while selected modes are validated through time-domain extraction when necessary.
@@
Table 1: QNMs of the $\ell=2$, $n=0$ axial gravitational perturbations for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
@@
For $\ell=2$, the time-domain integration yields
$\omega=0.369359-0.0856186i$
for the least damped mode, while the sixth-order WKB method with Padé approximants gives
@@
amounts to about $4\%$ for the damping rate and more than $1\%$ for the real oscillation frequency. In other words, for the lowest multipole $\ell=2$, the physical effect is at least one order of magnitude larger than the numerical uncertainty of the WKB approximation (see Figure 3). Moreover, as $\ell$ increases, the WKB error decreases significantly, becoming two or more orders of magnitude smaller than the effect itself (see Figure  4). Therefore, we conclude that the WKB method provides a reliable tool in this context. We can see that once the quantum correction is included, the real oscillation frequency and damping rate are decreased in comparison with their classical (Schwarzschild) limits.
Figure 3: Time-domain profile for $\ell=2$, $l_{cr}=1.137$. The Prony method gives $\omega=0.369359-0.0856186i$ for the fundamental mode $n=0$.
Figure 4: Time-domain profile for $\ell=3$, $l_{cr}=1.137$. The Prony method gives $\omega=0.369373-0.0856553i$ for the fundamental mode $n=0$.
@@ V Conclusions
In this work, we have analyzed the gravitational quasinormal spectrum of the Dymnikova black hole, a regular spacetime that can be interpreted either as a phenomenological model with a de Sitter core or as an effective geometry emerging from Asymptotically Safe gravity. Using the WKB method with Padé improvements, supported by time-domain integration, we have obtained the dominant quasinormal frequencies for axial gravitational perturbations. Our results show that the quantum correction parameter $l_{\rm cr}$ systematically modifies the spectrum: both the real oscillation frequencies and the damping rates decrease when the corrections are switched on, so that the ringdown signal becomes longer-lived compared to the Schwarzschild case. In the limit of large $l_{\rm cr}$, the Schwarzschild spectrum is smoothly recovered, confirming the consistency of the model.
It is important to emphasize that the present study has been restricted to the fundamental mode and the first overtone for which the WKB approach is reliable. For higher overtones, however, the situation is expected to be qualitatively different. Since overtones are strongly localized near the event horizon, they should be especially sensitive to the local geometric deformations induced by quantum corrections Konoplya and Zhidenko (2024c); Konoplya et al. (2022). To investigate this regime, one would need to resort to more precise techniques such as the Frobenius (Leaver) method Leaver (1985), which provides convergent series solutions and can capture the full spectrum beyond the eikonal approximation. At the same time, the metric, and consequently the master wave equation, does not have a rational form, so one has to transform or approximate the metric by some rational function for accurate calculations of higher overtones. Having in mind that the fundamental mode and possibly the first overtone are the only realistic candidates for detection in the near future experiments, we leave this analysis, as well as the inclusion of polar perturbations and possible extensions to higher dimensions, for future work.

### v2

@@ I Introduction
One of the most powerful probes of black-hole physics comes from the study of their quasinormal modes (QNMs) Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011); Bolokhov and Skvortsova (2025a), which govern the damped oscillations of spacetime following a perturbation. These modes encode key information about the underlying geometry and dominate the ringdown phase of gravitational-wave signals detected by LIGO, Virgo, and KAGRA Abbott et al. (2016a, 2017, 2020, b). The spectrum of gravitational perturbations provides a unique fingerprint of the compact object, enabling tests of general relativity in the strong-field regime and offering the possibility of constraining or ruling out alternative models of gravity and exotic compact objects. For this reason, accurate determination of gravitational QNMs has become a central task in both theoretical and observational astrophysics.
@@
Among the various proposals, the solution introduced by Dymnikova Dymnikova (1992) stands out as an elegant example. It describes a static, spherically symmetric, asymptotically flat geometry in which the central singularity is replaced by a de Sitter core, leading to finite curvature invariants everywhere. While originally formulated as a phenomenological construction, the same functional form has more recently reappeared in the framework of Asymptotic Safety, where it can be derived from renormalization-group improvements tied to curvature invariants Platania (2019). Thus, the Dymnikova spacetime provides not only one of the earliest regular black-hole models, but also a concrete realization of quantum-gravity inspired corrections. This dual interpretation makes it a natural and timely candidate for detailed studies of gravitational perturbations and their observational consequences.
@@ II The Dymnikova Black Hole Geometry
The quest for singularity-free black-hole solutions has long motivated the construction of effective geometries that remain regular throughout spacetime. A particularly influential example was proposed by Dymnikova Dymnikova (1992), who suggested replacing the central singularity of the Schwarzschild solution with a de Sitter core while preserving asymptotic flatness. This is achieved through a stress–energy distribution that formally violates some of the classical energy conditions but can be regarded as an effective description of quantum backreaction. The outcome is a static, spherically symmetric, and asymptotically flat spacetime in which all curvature invariants remain finite.
@@
describing a regular de Sitter interior, while at large $r$ one recovers the Schwarzschild behavior $f(r)\approx 1-2M/r$. The spacetime thus interpolates smoothly between two familiar limits while remaining free of singularities.
Although originally introduced phenomenologically, essentially the same functional form was later derived within the framework of Asymptotic Safety. In this approach, the Schwarzschild lapse function
@@
The Dymnikova solution thus enjoys a dual origin: historically as a phenomenological model of singularity resolution and, more recently, as a concrete realization of renormalization-group improvements in Asymptotic Safety. This dual perspective underscores its relevance as a minimal yet robust model of a regular black hole. Furthermore, higher-dimensional generalizations of the Dymnikova geometry have been shown to arise naturally in theories with higher-curvature corrections Konoplya and Zhidenko (2024a), further highlighting its role as a testing ground for semiclassical and quantum-gravity inspired modifications of black-hole physics.
@@ III Gravitational perturbations
Studying gravitational perturbations in spacetimes that incorporate quantum corrections is more subtle than in the classical case. The difficulty arises because such geometries, including the Dymnikova black hole, are not obtained as exact solutions of Einstein’s equations with a specified matter source, but rather emerge from effective constructions motivated by quantum-gravity considerations. As a result, a straightforward derivation of perturbation equations is not possible. A practical resolution was suggested in the loop-inspired framework of Ashtekar, Olmedo, and Singh Ashtekar et al. (2018a, b), where the background geometry can be reinterpreted within Einstein’s theory if one introduces an effective anisotropic fluid stress tensor that captures the relevant quantum corrections. This reformulation allows the machinery of black-hole perturbation theory to be applied in a consistent manner.
Following this idea, Bouhmadi-López and collaborators Bouhmadi-López et al. (2020); Konoplya and Stashko (2025a) analyzed axial perturbations under the assumption that fluctuations aligned with the anisotropy do not contribute in the axial sector. This simplification is reminiscent of the treatment in Bronnikov et al. (2012), where the decoupling of scalar and axial gravitational modes was exploited in scalar–tensor models (see also Chen and Chen (2019)). Although such assumptions necessarily restrict the analysis, they have been widely adopted in related contexts and are generally expected to provide reliable results for small deviations from Schwarzschild geometry Konoplya and Stashko (2025b, a); Dubinsky (2025a); Bolokhov and Skvortsova (2025c). For Dymnikova black holes as solutions of the Einstein equations with the energy-momentum tensor corresponding to some density profile, the perturbation equations were considered in Dymnikova and Galaktionov (2005).
@@
Figures 1 and 2 show the effective potentials for axial modes in the Dymnikova background. When the quantum parameter $l_{\rm cr}$ is very small (e.g., $l_{\rm cr}=0.01M$), the potential nearly coincides with the Schwarzschild one except in the immediate vicinity of the horizon. The deformation is strongly localized, and the potential rapidly merges with the Schwarzschild profile at moderate radii. At the same time, the location of the event horizon shifts inward as $l_{\rm cr}$ grows, leading to a decrease in the horizon radius compared to the Schwarzschild value. These features indicate that the quantum-inspired modifications are concentrated in the near-horizon region, leaving the asymptotic structure essentially unaffected.
@@ IV Quasinormal modes
The WKB–Padé method is most reliable for moderate to large multipoles and low overtones (heuristically $n\ll\ell$ for single-barrier potentials). Tables I–VIII summarize the results obtained with 6th and higher-order WKB methods combined with Padé resummation, while selected modes are validated through time-domain extraction when necessary.
@@
Table 1: QNMs of the $\ell=2$, $n=0$ test axial gravitational perturbations for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
@@
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.8091780.094164i0.809178-0.094164i 0.8091780.094164i0.809178-0.094164i $0\%$
$0.5$ 0.8091780.094164i0.809178-0.094164i 0.8091780.094164i0.809178-0.094164i $0\%$
$0.7$ 0.8091850.094154i0.809185-0.094154i 0.8091870.094169i0.809187-0.094169i $0.00193\%$
$0.9$ 0.8088940.094354i0.808894-0.094354i 0.8089150.094290i0.808915-0.094290i $0.00821\%$
$1.13$ 0.8066840.093368i0.806684-0.093368i 0.8067460.093185i0.806746-0.093185i $0.0239\%$
$1.137$ 0.8067090.093176i0.806709-0.093176i 0.8066340.093089i0.806634-0.093089i $0.0142\%$
Table 5: QNMs of the $\ell=4$, $n=0$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.7966310.284334i0.796631-0.284334i 0.7966320.284334i0.796632-0.284334i $0.00003\%$
$0.5$ 0.7966310.284334i0.796631-0.284334i 0.7966310.284334i0.796631-0.284334i $0.00002\%$
$0.7$ 0.7966370.284213i0.796637-0.284213i 0.7967250.284362i0.796725-0.284362i $0.0204\%$
$0.9$ 0.7947300.285220i0.794730-0.285220i 0.7949160.284828i0.794916-0.284828i $0.0515\%$
$1.13$ 0.7855380.280464i0.785538-0.280464i 0.7851650.280327i0.785165-0.280327i $0.0476\%$
$1.137$ 0.7849540.280114i0.784954-0.280114i 0.7846720.280018i0.784672-0.280018i $0.0357\%$
Table 6: QNMs of the $\ell=4$, $n=1$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l_{cr}$ WKB-6 $m=3$ WKB-7 $m=4$ difference
$0.3$ 0.7727100.479903i0.772710-0.479903i 0.7727120.479908i0.772712-0.479908i $0.00060\%$
$0.5$ 0.7727100.479903i0.772710-0.479903i 0.7727080.479906i0.772708-0.479906i $0.00039\%$
$0.7$ 0.7724140.479325i0.772414-0.479325i 0.7731600.479874i0.773160-0.479874i $0.102\%$
$0.9$ 0.7648890.481360i0.764889-0.481360i 0.7653380.480293i0.765338-0.480293i $0.128\%$
$1.13$ 0.7445440.468344i0.744544-0.468344i 0.7410810.468153i0.741081-0.468153i $0.394\%$
$1.137$ 0.7431630.467698i0.743163-0.467698i 0.7399200.467587i0.739920-0.467587i $0.370\%$
Table 7: QNMs of the $\ell=4$, $n=2$ for the Dymnikova black hole $M=1$ calculated using the WKB formula at different orders and Padé approximants.
$l$ WKB-6 $m=3$ WKB-7 $m=4$ difference
22 0.3689990.085650i0.368999-0.085650i 0.3689560.085529i0.368956-0.085529i $0.0341\%$
33 0.5959270.090979i0.595927-0.090979i 0.5961490.091043i0.596149-0.091043i $0.0384\%$
44 0.8067090.093176i0.806709-0.093176i 0.8066340.093089i0.806634-0.093089i $0.0142\%$
55 1.0102920.094095i1.010292-0.094095i 1.0102460.094067i1.010246-0.094067i $0.00535\%$
66 1.2103410.094617i1.210341-0.094617i 1.2103100.094606i1.210310-0.094606i $0.00266\%$
77 1.4083140.094934i1.408314-0.094934i 1.4082950.094930i1.408295-0.094930i $0.00143\%$
88 1.6049650.095140i1.604965-0.095140i 1.6049520.095139i1.604952-0.095139i $0.00080\%$
99 1.8007180.095280i1.800718-0.095280i 1.8007090.095281i1.800709-0.095281i $0.00047\%$
$10$ 1.9958340.095379i1.995834-0.095379i 1.9958290.095380i1.995829-0.095380i $0.00028\%$
$15$ 2.9662220.095605i2.966222-0.095605i 2.9662210.095606i2.966221-0.095606i $0.00003\%$
2020 3.9326290.095679i3.932629-0.095679i 3.9326290.095679i3.932629-0.095679i $5.4\times 10^{-6}\%$
$25$ 4.8974090.095711i4.897409-0.095711i 4.8974090.095711i4.897409-0.095711i $1.2\times 10^{-6}\%$
$30$ 5.8613660.095728i5.861366-0.095728i 5.8613660.095728i5.861366-0.095728i $0\%$
$35$ 6.8248490.095738i6.824849-0.095738i 6.8248490.095738i6.824849-0.095738i $0\%$
$40$ 7.7880340.095744i7.788034-0.095744i 7.7880340.095744i7.788034-0.095744i $0\%$
$45$ 8.7510200.095748i8.751020-0.095748i 8.7510200.095748i8.751020-0.095748i $0\%$
$50$ 9.7138650.095751i9.713865-0.095751i 9.7138650.095751i9.713865-0.095751i $0\%$
$75$ 14.5269640.095758i14.526964-0.095758i 14.5269640.095758i14.526964-0.095758i $0\%$
$100$ 19.3392140.095761i19.339214-0.095761i 19.3392140.095761i19.339214-0.095761i $0\%$
$150$ 28.9628610.095762i28.962861-0.095762i 28.9628610.095762i28.962861-0.095762i $0\%$
$200$ 38.5860780.095763i38.586078-0.095763i 38.5860780.095763i38.586078-0.095763i $0\%$
Table 8: Quasinormal modes for various $\ell$, calculated using the WKB formula at different orders and Pade approximants. Here we have $n=0$, $M=1$, $l_{cr}=1.137$.
@@
When using high accuracy time-domain integration, for $\ell=2$, the time-domain integration yields
$\omega=0.369372-0.0856535i$
for the least damped mode. This value is extracted by the Prony method with all accuracy of all 6 digits when fitting at the intervals $t=(150,200)$ and $t=(200,250)$. Thus, there is no point in extending computations to later times. The sixth-order WKB method with Padé approximants gives
@@
amounts to about $4\%$ for the damping rate and more than $1\%$ for the real oscillation frequency. In other words, for the lowest multipole $\ell=2$, the physical effect is at least one order of magnitude larger than the numerical uncertainty of the WKB approximation (see Figure 3). Moreover, as $\ell$ increases, the WKB error decreases significantly, becoming two or more orders of magnitude smaller than the effect itself (see Figure 4). Therefore, we conclude that the WKB method provides a reliable tool in this context. We can see that once the quantum correction is included, the real oscillation frequency and damping rate are decreased in comparison with their classical (Schwarzschild) limits.
Figure 3: Time-domain profile for $\ell=2$, $l_{cr}=1.137$. The Prony method gives $\omega=0.369372-0.0856535i$ for the fundamental mode $n=0$.
Figure 4: Time-domain profile for $\ell=3$, $l_{cr}=1.137$. The Prony method gives $\omega=0.59618-0.091002i$ for the fundamental mode $n=0$.
@@
Correspondence between QNMs and null geodesics. In the eikonal (geometric–optics) regime, when the multipole number $\ell$ is large, the dynamics of perturbations around a black hole is closely related to the properties of unstable circular null geodesics. In this limit, the wavefronts of the perturbing field propagate along null rays, and the effective potential governing quasinormal oscillations develops a sharp peak near the photon sphere. The real part of the quasinormal frequency is then determined by the angular velocity $\Omega_{c}$ of the null particle at this orbit, while the imaginary part is governed by the Lyapunov exponent $\lambda$, which characterizes the instability timescale of the corresponding geodesic trajectory Cardoso et al. (2009). Consequently, the fundamental quasinormal frequencies can be approximated as
$\omega_{\ell n}\simeq\Omega_{c}\,\ell-i\left(n+\frac{1}{2}\right)\lambda,$ (20)
where $n$ is the overtone number. This correspondence holds for a broad class of black holes with single–peaked effective potentials, providing an intuitive geometrical interpretation of the quasinormal spectrum in terms of the motion of photons on unstable circular orbits. Nevertheless, there are a number of deviations from this relation, which are described in Khanna and Price (2017); Konoplya and Stuchlík (2017); Bolokhov (2024d); Konoplya (2023) and are usually related to gravitational or other non-minimally coupled perturbations. They may signal the breakdown of the eikonal approximation. Here, in Table VIII we study the high $\ell$ regime of QNMs for the near-extreme Dymnikova black hole. One can easily compute the proper characteristics of the null geodesics for the Dymnikov black hole and compare them with the numerical data presented in Table VIII. We conclude that the correspondence holds for gravitational perturbations in our case.
@@ V Conclusions
In this work, we have analyzed the gravitational quasinormal spectrum of the Dymnikova black hole, a regular spacetime which can be interpreted either as a phenomenological model with a de Sitter core or as an effective geometry emerging from Asymptotically Safe gravity. Using the WKB method with Padé improvements, supported by time-domain integration, we have obtained the dominant quasinormal frequencies for axial gravitational perturbations. Our results show that the quantum correction parameter $l_{\rm cr}$ systematically modifies the spectrum: both the real oscillation frequencies and the damping rates decrease when the corrections are switched on, so that the ringdown signal becomes longer-lived compared to the Schwarzschild case. In the limit of large $l_{\rm cr}$, the Schwarzschild spectrum is smoothly recovered, confirming the consistency of the model. We also analyzed the regime of large multipole numbers and confirmed the correspondence between null geodesics and eikonal QNMs.
It is important to emphasize that the present study has been mainly restricted to the fundamental mode and the first overtone for which the WKB approach is reliable. For higher overtones, however, the situation is expected to be qualitatively different. Since overtones are strongly localized near the event horizon, they should be especially sensitive to the local geometric deformations induced by quantum corrections Konoplya and Zhidenko (2024c); Konoplya et al. (2022). To investigate this regime, one would need to resort to more precise techniques such as the Frobenius (Leaver) method Leaver (1985), which provides convergent series solutions and can capture the full spectrum beyond the eikonal approximation. At the same time, the metric, and consequently the master wave equation, does not have a rational form, so one has to transform or approximate the metric by some rational function for accurate calculations of higher overtones. Having in mind that the fundamental mode and possibly the first overtone are the only realistic candidates for detection in the near future experiments, we leave this analysis, as well as the inclusion of polar perturbations and possible extensions to higher dimensions, for future work.